A 75-kg man pushes a 10-kg cart along a 25-m corridor with a 1500-N force at an angle of 30O with the horizontal at a µ of 0.45.Find the work done.


1
Expert's answer
2021-02-23T10:02:38-0500

The work done by a man

"W=Fd\cos\theta""=\rm 1500\: N\times 25\: m\times \cos 30^{\circ}=32476\: J=32.5\: kJ"